OpenStudy (anonymous):

the laplace tranform of -tcosx=-cosx//s^2 am i right???

OpenStudy (amistre64):

doesnt look right to me, might be your choice of notation tho

OpenStudy (amistre64):

do we agree that: L(y') = sL(y) - yo

OpenStudy (anonymous):

my question is different

OpenStudy (amistre64):

is it about bubble gum? because it looks to be about a laplace transform to me ... can you clarify it any?

OpenStudy (amistre64):

here is how Im reading you question since you do not wish to clarify your responses using the product rule of derivatives L(tg' + g) = sL(tg) - 0go, by linearity of the laplace we get L(tg') + L(g) = sL(tg) let g = -cos(x) g' = sin(x) sL(-tcos(x)) = L(tsin(x)) + L(-cos(x)) s^2 L(-tcos(x)) = sL(tsin(x)) + sL(-cos(x)) using the same construct we can determine the sL(tsin(x)) ---------------- sL(tsin(x)) = L(tcos(x)) + L(sin(x)) s^2 L(-tcos(x)) = L(tcos(x)) + L(sin(x)) + sL(-cos(x)) s^2 L(-tcos(x)) + L(-tcos(x)) = L(sin(x)) + sL(-cos(x)) (s^2+1) L(-tcos(x)) = L(sin(x)) + sL(-cos(x)) L(-tcos(x)) = [L(sin(x)) + sL(-cos(x))]/(s^2+1)

OpenStudy (amistre64):

your questions is in x and t ... is that a typo? or on purpose?

OpenStudy (amistre64):

assuming cos(x) is not a typo, then its simply a constant with respect to t

OpenStudy (anonymous):

sory for bieng late rep but i have some personal problems so thats y.

OpenStudy (anonymous):

yes cosx is treat constant.

OpenStudy (anonymous):

then the laplace is -cosx/s^2 is right?

OpenStudy (math&ing001):

Wolfram says it's right.

OpenStudy (anonymous):

wow its amazing.

OpenStudy (amistre64):

late is better then never :)

OpenStudy (math&ing001):

You can use the tool from here http://www.wolframalpha.com/widgets/view.jsp?id=3272010f63d3145699ca78bbe0db05a7

OpenStudy (anonymous):

thank u nice to meet u.